If sinA – sinB = c and cosA – cosB = d then cosec2 [(A + B)/2] is equal to

Grade:12th pass

1 Answers

Vishal Kumar
5 years ago
Sin(A) - sin( B) = 2 sin ( A-B /2) cos (A +B/2) = c....1
Cos(A)- cos(B)= 2 sin ( B - A/2)sin ( A+B/2)= d....2
Dividing equation 1 and 2
Cot ( A+B/2) = -c/d 
Now as we know that
Cot2x + cosec2x =11
So now calculate cosec2( A+B/2) = 1-( c/d)2